OpenStudy (anonymous):

6. Cleo and Clare are looking from their balcony to a swimming pool below that is located 15 m horizontally from the bottom of their building. They estimate the balcony is 45 m high and wonder how fast they would have to jump horizontally to succeed in reaching the pool. What calculations would you show to help them determine the answer? Evaluate the practicality of their being able to succeed at jumping into the pool.

OpenStudy (anonymous):

@Tumblewolf please help

OpenStudy (tumblewolf):

In this problem, you do vertical first, to determine how long it would take to fall from the balcony to the ground. You know that the initial displacement is 45 m, the final displacement is zero (ground level), and the initial vertical velocity is zero (they jump horizontally). So, you have: Dfinal = Dinitial + Vinitial*t + 0.5*a*t^2 0 = 45 + 0 - 0.5*9.81*t^2 t^2 = 45/(0.5*9.81) t = sqrt(45/(0.5*9.81) Can you do the math from here?

OpenStudy (tumblewolf):

If not then in easier terms we have s=ut+1/2at^2 45=0+1/2gt^2 45=4.9t^2 9.184=t^2 t = 3.03 seconds from balcony to the pool